P(x)=x^3-9x^2+11x-3
Compute P(1) and P(-1)

Answer by stanbon(75887)

P(x)=x^3-9x^2+11x-3
Compute
P(1) = 1-9+11-3 = 0
--------------------------
P(-1) = -1-9-11-3 = -24
============================
Cheers,
Stan H.